# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v3

@@ C.4 Proof of Lemma 9
To prove Lemma 9, we relate the magnitude of error in trace distance to that in the parameter $\bm{\theta}$. For this purpose, we begin by proving several lemmas. Since any compact Lie group $G$ is isomorphic to a closed linear group, we assume $G$ is a compact linear Lie group and adopt its Schatten norm in the following arguments.
@@
$\|\mathcal{U}_{g}(\rho)-\mathcal{U}_{\hat{g}}(\rho)\|_{1}<\delta_{*}$ $\implies\exists\bm{\varphi}\,\text{ s.t. }\begin{cases}&\mathcal{U}_{\bm{\varphi}}\circ\mathcal{U}_{\hat{g}}(\rho)=\mathcal{U}_{g}(\rho)\\ &\|\bm{\varphi}\|\leq c\|\mathcal{U}_{g}(\rho)-\mathcal{U}_{\hat{g}}(\rho)\|_{1}\end{cases},$ (363)
where $\mathcal{U}_{\bm{\varphi}}(\cdot)\coloneqq e^{\mathrm{i}\bm{\varphi}\cdot\bm{Y}}(\cdot)e^{-\mathrm{i}\bm{\varphi}\cdot\bm{Y}}$ with $\bm{\varphi}\in\mathbb{R}^{n}$ and $\bm{Y}=\{Y_{i}\}_{i=1}^{n}$ is a basis of the linear span of $\{X_{\mu}\}_{\mu=1}^{\dim G}$, where $X_{\mu}$ is defined in Eq. (4).
@@
and $c$ to be arbitrary positive number, Eq. (384) holds for $\bm{\varphi}=\bm{0}$. We remark that this essentially includes the case where $G$ is a finite group.
@@
Fix arbitrary $g,\hat{g}\in G$ satisfying $\|\mathcal{U}_{g}(\rho)-\mathcal{U}_{\hat{g}}(\rho)\|_{1}<\delta$, and find appropriate $\bm{\varphi}$ satisfying Eq. (384). By the definition of $\delta$, we can take $h\in\mathrm{Sym}_{G}(\rho)$ such that $\|U(g^{-1}\hat{g}h)-I\|_{\infty}\leq 1-e^{-1/8}<1$. Thus we can define
@@
$\|\bm{\theta}\|\leq\frac{1}{lm}\|U(\hat{g})\rho U(\hat{g})^{\dagger}-U(g)\rho U(g)^{\dagger}\|_{1},$ (377)
@@
Since we can take $\bm{\gamma}$ satisfying $\bm{\varphi}=\|\bm{\varphi}\|\bm{\gamma}$ and $\|\bm{\gamma}\|=1$, we have
$\|B\|_{\infty}=\left\|\sum_{k}\varphi^{k}Y_{k}\right\|_{\infty}=\|\bm{\varphi}\|\left\|\sum_{k}\varphi^{k}Y_{k}\right\|_{\infty}\geq l\|\bm{\varphi}\|.$ (382)
@@
$\|\bm{\varphi}\|\leq\frac{\|B\|_{\infty}}{l}\leq\frac{\|A_{1}\|_{\infty}}{2l}\leq\frac{1}{lm}\|U(\hat{g})\psi U(\hat{g})^{\dagger}-U(g)\psi U(g)^{\dagger}\|_{\infty}\leq\frac{1}{lm}\|U(\hat{g})\psi U(\hat{g})^{\dagger}-U(g)\psi U(g)^{\dagger}\|_{1}.$ (383)
@@
$\|\mathcal{U}_{g}(\rho)-\mathcal{U}_{\hat{g}}(\rho)\|_{1}<\delta_{*}$ $\implies\exists\bm{\theta}\,\text{ s.t. }\begin{cases}&\mathcal{U}_{\bm{\theta}}\circ\mathcal{U}_{\hat{g}}(\rho)=\mathcal{U}_{g}(\rho)\\ &\|\bm{\theta}\|\leq c\|\mathcal{U}_{g}(\rho)-\mathcal{U}_{\hat{g}}(\rho)\|_{1}\end{cases},$ (384)
where $\mathcal{U}_{\bm{\theta}}(\cdot)\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}}(\cdot)e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}}$ for $\bm{\theta}\in\mathbb{R}^{\dim G}$, where $\bm{X}=\{X_{\mu}\}_{\mu=1}^{\dim G}$ is defined in Eq. (4).
@@
Since $\bm{Y}=\{Y_{i}\}_{i=1}^{n}$ in Lemma 39 is a basis of the linear span of $\{X_{\mu}\}_{\mu=1}^{\dim G}$, there is an $n\times\dim G$ matrix $L$ such that $L\bm{X}=\bm{Y}$. By using the transpose $L^{\top}$ of $L$, we define $\bm{\theta}\in\mathbb{R}^{\dim G}$ by $\bm{\theta}\coloneqq L^{\top}\bm{\varphi}$, which satisfies $\bm{\theta}\cdot\bm{X}=\bm{\varphi}\cdot\bm{Y}$. Note that $\|\bm{\theta}\|\leq\|L^{\top}\|_{2}\|\bm{\varphi}\|$ holds, where $\|L^{\top}\|_{2}$ denotes the matrix norm induced from the 2-norm for vectors, which is finite. Therefore, Lemma 40 holds by replacing the constant $c$ in Lemma 39 to $c\|L^{\top}\|_{2}$. ∎
@@
by performing measurements on $n=\lceil N^{1-\epsilon}\rceil$ copies of the system in a state $\mathcal{U}_{g}(\rho)^{\otimes n}$, where we have used $nN^{-1+2\epsilon}\geq N^{1-\epsilon}N^{-1+2\epsilon}=N^{\epsilon}$. From the estimated state $\hat{\rho}_{n}$, we pick up any $\hat{g}\in G$ satisfying $\|\mathcal{U}_{\hat{g}}(\rho)-\hat{\rho}_{n}\|_{1}<N^{-1/2+\epsilon}/(2c)$ as an estimate of the true value $g\in G$. If there is no $\hat{g}$ satisfying $\|\mathcal{U}_{\hat{g}}(\rho)-\hat{\rho}_{n}\|_{1}<N^{-1/2+\epsilon}/(2c)$, which may happen with an exponentially small probability, we arbitrarily select $\hat{g}$ as an element in $G$, such as $e\in G$. Whenever $\|\hat{\rho}_{n}-\mathcal{U}_{g}(\rho)\|_{1}<N^{-1/2+\epsilon}/(2c)$ holds, the estimated value $\hat{g}\in G$ satisfies $\|\mathcal{U}_{\hat{g}}(\rho)-\mathcal{U}_{g}(\rho)\|_{1}<N^{-1/2+\epsilon}/c$ due to the triangle inequality. In this case, for all sufficiently large $N$ such that $N^{-1/2+\epsilon}/c<\delta_{*}$, we get $\|\bm{\theta}\|<N^{-1/2+\epsilon}$. Therefore, the success probability is bounded as
@@ Appendix D Proof of Lemma 10
[…]
$\delta_{N}(g,\bm{u})\coloneqq T\left(\mathcal{E}^{(g)}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\left(\psi_{g}\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}\left(\phi_{g}\right)^{\otimes\lfloor rN\rfloor}\right)$ (390)
satisfies $\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon}}\delta_{N}(g,\bm{u})=0$ for $\epsilon\in(0,1/9)$.
@@
Let $G=\bigsqcup_{i=0}^{k}G_{i}$ be the decomposition of a compact Lie group $G$ into its connected components. We arbitrarily fix $g_{i}\in G_{i}$ for $i=0,\cdots,k$. Applying Lemma 8 to $\psi_{g_{i}}\coloneqq\mathcal{U}_{g_{i}}$ and $\phi_{g_{i}}\coloneqq\mathcal{U}^{\prime}_{g_{i}}(\phi)$ for each $g_{i}$, we find that there exists a sequence of quantum channels $\{\mathcal{E}_{N}^{(g_{i})}\}_{N}$ such that
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi_{g_{i}})^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\phi_{g_{i}})^{\otimes\lfloor rN\rfloor}\right)=0$ (391)
@@
For a general $g\in G$, if $g_{i}$ belongs to a connected component $g\in G_{i}$, we define a channel $\mathcal{E}_{N}^{(g)}$ by using $\mathcal{E}_{N}^{(g_{i})}$ as follows: Let $g^{\prime}\in G_{0}$ be an element in the identity component satisfying $g=hg_{i}$. From Eq. (129), there exists an invertible $\dim G\times\dim G$ real matrix $V(h)$ satisfying
$U(h)^{\dagger}X_{\mu}U(h)=\sum_{\nu=1}^{\dim G}V(h)_{\nu\mu}X_{\nu}.$ (392)
We define $\mathcal{E}_{N}^{(g)}\coloneqq\left(\mathcal{U}^{\prime}_{h}\right)^{\otimes\lfloor rN\rfloor}\circ\mathcal{E}_{N}^{(g_{i})}\circ\left(\mathcal{U}_{h}^{-1}\right)^{\otimes N}$. Since $U(h)^{\dagger}(\bm{u}\cdot\bm{X})U(h)=\bm{u}\cdot V(h)^{\top}\bm{X}=\bm{v}\cdot\bm{X}$, where $\bm{v}\coloneqq\bm{u}V(h)^{\top}$, we have
$\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\left(\psi_{g}\right)^{\otimes N}\right)$ $=\left(\mathcal{U}_{h}\right)^{\otimes\lfloor rN\rfloor}\circ\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{h}^{-1}\circ\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\circ\mathcal{U}_{h}\left(\psi_{g_{i}}\right)\right)^{\otimes N}$ (393)
$=\left(\mathcal{U}^{\prime}_{h}\right)^{\otimes\lfloor rN\rfloor}\circ\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{\frac{\bm{v}}{\sqrt{N}}}\left(\psi_{g_{i}}\right)\right).$ (394)
@@
$T\left(\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi_{g})^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\phi_{g})^{\otimes\lfloor rN\rfloor}\right)$ $=T\left(\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{\frac{\bm{v}}{\sqrt{N}}}\left(\psi_{g_{i}}\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{v}}{\sqrt{N}}}\left(\phi_{g_{i}}\right)^{\otimes\lfloor rN\rfloor}\right),$ (395)
where we have used $U^{\prime}(h)^{\dagger}(\bm{u}\cdot\bm{X}^{\prime})U^{\prime}(h)=\bm{v}\cdot\bm{X}^{\prime}$. Note that the induced norm $\|V(h)\|_{\infty}\coloneqq\max_{\bm{u}\neq\bm{0}}\|V(h)\bm{u}\|/\|\bm{u}\|$ is finite for any $h\in G_{0}$. Since $\|V(h)\|_{\infty}$ and continuous in $h\in G_{0}$ and $G_{0}$ is compact, we find $c\coloneqq\sup_{h\in G_{0}}\|V(h)\|_{\infty}$ is also finite. By using $\|\bm{v}\|\leq c\|\bm{u}\|$, for any $g\in G$ and $\|\bm{u}\|<N^{\epsilon}$, we get
$\delta_{N}(g,\bm{u})$ $\coloneqq T\left(\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi_{g})^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\phi_{g})^{\otimes\lfloor rN\rfloor}\right)$ (396)
$\leq\max_{i\in\{0,\cdots,k\}}\sup_{\|\bm{v}\|<cN^{\epsilon}}T\left(\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{\frac{\bm{v}}{\sqrt{N}}}\left(\psi_{g_{i}}\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{v}}{\sqrt{N}}}\left(\phi_{g_{i}}\right)^{\otimes\lfloor rN\rfloor}\right).$ (397)
@@
$\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon}}\delta_{N}(g,\bm{u})$ $\leq\lim_{N\to\infty}\max_{i\in\{0,\cdots,k\}}\sup_{\|\bm{v}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{\frac{\bm{v}}{\sqrt{N}}}\left(\psi_{g_{i}}\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{v}}{\sqrt{N}}}\left(\phi_{g_{i}}\right)^{\otimes\lfloor rN\rfloor}\right)$ (398)
$=\max_{i\in\{0,\cdots,k\}}\lim_{N\to\infty}\sup_{\|\bm{v}\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}_{N}^{(g_{i})}\left(\mathcal{U}_{\frac{\bm{v}}{\sqrt{N}}}\left(\psi_{g_{i}}\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{v}}{\sqrt{N}}}\left(\phi_{g_{i}}\right)^{\otimes\lfloor rN\rfloor}\right)$ (399)

### v4

@@ D.2.4 Proof of Lemma 26
To prove Lemma 26, we relate the magnitude of error in trace distance to that in the parameter $\theta$. For this purpose, we begin by proving several lemmas. Since any compact Lie group $G$ is isomorphic to a closed linear group, we assume $G$ is a compact linear Lie group and adopt its Schatten norm in the following arguments.
@@
$\|\mathcal{U}_{g}(\rho)-\mathcal{U}_{\hat{g}}(\rho)\|_{1}<\delta_{*}$ $\implies\exists\varphi\,\text{ s.t. }\begin{cases}&\mathcal{U}_{\varphi}\circ\mathcal{U}_{\hat{g}}(\rho)=\mathcal{U}_{g}(\rho)\\ &\|\varphi\|\leq c\|\mathcal{U}_{g}(\rho)-\mathcal{U}_{\hat{g}}(\rho)\|_{1}\end{cases},$ (365)
where $\mathcal{U}_{\varphi}(\cdot)\coloneqq e^{\mathrm{i}\sum_{k=1}^{n}\varphi^{k}Y_{k}}(\cdot)e^{-\mathrm{i}\sum_{k=1}^{n}\varphi^{k}Y_{k}}$ with $\varphi\in\mathbb{R}^{n}$ and $\{Y_{k}\}_{k=1}^{n}$ is a basis of the linear span of $\{X_{\mu}\}_{\mu=1}^{\dim G}$, where $X_{\mu}$ is defined in Eq. (6).
@@
and $c$ to be arbitrary positive number, Eq. (365) holds for $\varphi=0$. We remark that this essentially includes the case where $G$ is a finite group.
@@
Fix arbitrary $g,\hat{g}\in G$ satisfying $\|\mathcal{U}_{g}(\rho)-\mathcal{U}_{\hat{g}}(\rho)\|_{1}<\delta$, and find appropriate $\varphi$ satisfying Eq. (365). By the definition of $\delta$, we can take $h\in\mathrm{Sym}_{G}(\rho)$ such that $\|U(g^{-1}\hat{g}h)-I\|_{\infty}\leq 1-e^{-1/8}<1$. Thus we can define
@@
$\|\varphi\|\leq\frac{1}{lm}\|U(\hat{g})\rho U(\hat{g})^{\dagger}-U(g)\rho U(g)^{\dagger}\|_{1},$ (379)
@@
Since we can take $\gamma$ satisfying $\varphi=\|\varphi\|\gamma$ and $\|\gamma\|=1$, we have
$\|B\|_{\infty}=\left\|\sum_{k}\varphi^{k}Y_{k}\right\|_{\infty}=\|\varphi\|\left\|\sum_{k}\varphi^{k}Y_{k}\right\|_{\infty}\geq l\|\varphi\|.$ (384)
@@
$\|\varphi\|\leq\frac{\|B\|_{\infty}}{l}\leq\frac{\|A_{1}\|_{\infty}}{2l}\leq\frac{1}{lm}\|U(\hat{g})\psi U(\hat{g})^{\dagger}-U(g)\psi U(g)^{\dagger}\|_{\infty}\leq\frac{1}{lm}\|U(\hat{g})\psi U(\hat{g})^{\dagger}-U(g)\psi U(g)^{\dagger}\|_{1}.$ (385)
@@
$\|\mathcal{U}_{g}(\rho)-\mathcal{U}_{\hat{g}}(\rho)\|_{1}<\delta_{*}$ $\implies\exists\theta\,\text{ s.t. }\begin{cases}&\mathcal{U}_{\theta}\circ\mathcal{U}_{\hat{g}}(\rho)=\mathcal{U}_{g}(\rho)\\ &\|\theta\|\leq c\|\mathcal{U}_{g}(\rho)-\mathcal{U}_{\hat{g}}(\rho)\|_{1}\end{cases},$ (386)
where $\mathcal{U}_{\theta}(\cdot)\coloneqq e^{\mathrm{i}\theta^{\mu}X_{\mu}}(\cdot)e^{-\mathrm{i}\theta^{\nu}X_{\nu}}$ for $\theta\in\mathbb{R}^{\dim G}$, where $X_{\mu}$ is defined in Eq. (6).
@@
Since $\{Y_{i}\}_{i=1}^{n}$ in Lemma 39 is a basis of the linear span of $\{X_{\mu}\}_{\mu=1}^{\dim G}$, there is an $n\times\dim G$ matrix $L$ such that $\sum_{\mu=1}^{\dim G}L_{i\mu}X_{\mu}=Y_{i}$. By using the transpose $L^{\top}$ of $L$, we define $\theta\in\mathbb{R}^{\dim G}$ by $\theta\coloneqq L^{\top}\varphi$, which satisfies $\theta^{\mu}X_{\mu}=\sum_{k}\varphi^{k}Y_{k}$. Note that $\|\theta\|\leq\|L^{\top}\|_{2}\|\varphi\|$ holds, where $\|L^{\top}\|_{2}$ denotes the matrix norm induced from the 2-norm for vectors, which is finite. Therefore, Lemma 40 holds by replacing the constant $c$ in Lemma 39 to $c\|L^{\top}\|_{2}$. ∎
@@
by performing measurements on $n=\lceil N^{1-\epsilon}\rceil$ copies of the system in a state $\mathcal{U}_{g}(\rho)^{\otimes n}$, where we have used $nN^{-1+2\epsilon}\geq N^{1-\epsilon}N^{-1+2\epsilon}=N^{\epsilon}$. From the estimated state $\hat{\rho}_{n}$, we pick up any $\hat{g}\in G$ satisfying $\|\mathcal{U}_{\hat{g}}(\rho)-\hat{\rho}_{n}\|_{1}<N^{-1/2+\epsilon}/(2c)$ as an estimate of the true value $g\in G$. If there is no $\hat{g}$ satisfying $\|\mathcal{U}_{\hat{g}}(\rho)-\hat{\rho}_{n}\|_{1}<N^{-1/2+\epsilon}/(2c)$, which may happen with an exponentially small probability, we arbitrarily select $\hat{g}$ as an element in $G$, such as $e\in G$. Whenever $\|\hat{\rho}_{n}-\mathcal{U}_{g}(\rho)\|_{1}<N^{-1/2+\epsilon}/(2c)$ holds, the estimated value $\hat{g}\in G$ satisfies $\|\mathcal{U}_{\hat{g}}(\rho)-\mathcal{U}_{g}(\rho)\|_{1}<N^{-1/2+\epsilon}/c$ due to the triangle inequality. In this case, for all sufficiently large $N$ such that $N^{-1/2+\epsilon}/c<\delta_{*}$, we get $\|\theta\|<N^{-1/2+\epsilon}$. Therefore, the success probability is bounded as
@@ D.2.5 Proof of Lemma 27
Let $U,U^{\prime}$ be (non-projective) unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Let $\psi\in\mathcal{P}(\mathcal{H})$ and $\phi\in\mathcal{P}(\mathcal{H}^{\prime})$ be pure states. If $r>0$ satisfies $\mathcal{Q}^{\mathcal{U}_{g}(\psi)}\geq r\mathcal{Q}^{\mathcal{U}_{g}^{\prime}(\phi)}$ for all $g\in G$, then, there exists a sequence of quantum channels $\{\mathcal{E}^{(g)}_{N}\}_{N}$ such that the conversion error
$\delta_{N}(g,u)\coloneqq T\left(\mathcal{E}^{(g)}_{N}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}\left(\mathcal{U}_{g}(\psi)\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}\left(\mathcal{U}_{g}^{\prime}(\phi)\right)^{\otimes\lfloor rN\rfloor}\right)$ (392)
satisfies $\lim_{N\to\infty}\sup_{g\in G}\sup_{\|u\|<N^{\epsilon}}\delta_{N}(g,u)=0$ for $\epsilon\in(0,1/9)$.
@@
Applying Lemma 25 to $\psi$ and $\phi$, we find that there exists a sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ such that
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{E}_{N}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\psi)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}(\phi)^{\otimes\lfloor rN\rfloor}\right)=0$ (393)
@@
For a general $g\in G$, we define a channel $\mathcal{E}_{N}^{(g)}$ by using $\mathcal{E}_{N}$ as follows: From Eq. (112), there exists an invertible $\dim G\times\dim G$ real matrix $V(g)$ satisfying
$U(g)^{\dagger}X_{\mu}U(g)=\sum_{\nu=1}^{\dim G}V(g)_{\nu\mu}X_{\nu}.$ (394)
We define $\mathcal{E}_{N}^{(g)}\coloneqq\left(\mathcal{U}^{\prime}_{g}\right)^{\otimes\lfloor rN\rfloor}\circ\mathcal{E}_{N}\circ\left(\mathcal{U}_{g}^{-1}\right)^{\otimes N}$. Since $U(g)^{\dagger}(u^{\mu}X_{\mu})U(g)=v^{\mu}X_{\mu}$, where $v\coloneqq uV(g)^{\top}$, for $\psi_{g}\coloneqq\mathcal{U}_{g}(\psi)$ we have
$\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}\left(\psi_{g}\right)^{\otimes N}\right)$ $=\left(\mathcal{U}_{g}\right)^{\otimes\lfloor rN\rfloor}\circ\mathcal{E}_{N}\left(\mathcal{U}_{g}^{-1}\circ\mathcal{U}_{\frac{u}{\sqrt{N}}}\circ\mathcal{U}_{g}\left(\psi\right)\right)^{\otimes N}$ (395)
$=\left(\mathcal{U}^{\prime}_{g}\right)^{\otimes\lfloor rN\rfloor}\circ\mathcal{E}_{N}\left(\mathcal{U}_{\frac{v}{\sqrt{N}}}\left(\psi\right)\right).$ (396)
@@
$T\left(\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\psi_{g})^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}(\phi_{g})^{\otimes\lfloor rN\rfloor}\right)$ $=T\left(\mathcal{E}_{N}\left(\mathcal{U}_{\frac{v}{\sqrt{N}}}\left(\psi\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{v}{\sqrt{N}}}\left(\phi\right)^{\otimes\lfloor rN\rfloor}\right),$ (397)
where we have used $U^{\prime}(h)^{\dagger}(u^{\mu}X_{\mu}^{\prime})U^{\prime}(h)=v^{\mu}X_{\mu}^{\prime}$. Note that the induced norm $\|V(g)\|_{\infty}\coloneqq\max_{u\neq 0}\|V(g)u\|/\|u\|$ is finite for any $g\in G$. Since $\|V(h)\|_{\infty}$ and continuous in $g\in G$ and $G$ is compact, we find $c\coloneqq\sup_{g\in G}\|V(g)\|_{\infty}$ is also finite. By using $\|v\|\leq c\|u\|$, for any $g\in G$ and $\|u\|<N^{\epsilon}$, we get
$\delta_{N}(g,u)$ $=T\left(\mathcal{E}_{N}^{(g)}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\psi_{g})^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}(\phi_{g})^{\otimes\lfloor rN\rfloor}\right)$ (398)
$\leq\sup_{\|v\|<cN^{\epsilon}}T\left(\mathcal{E}_{N}\left(\mathcal{U}_{\frac{v}{\sqrt{N}}}\left(\psi\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{v}{\sqrt{N}}}\left(\phi\right)^{\otimes\lfloor rN\rfloor}\right).$ (399)
@@
$\lim_{N\to\infty}\sup_{g\in G}\sup_{\|u\|<N^{\epsilon}}\delta_{N}(g,u)$ $\leq\lim_{N\to\infty}\sup_{\|v\|<N^{\epsilon^{\prime}}}T\left(\mathcal{E}_{N}\left(\mathcal{U}_{\frac{v}{\sqrt{N}}}\left(\psi\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{v}{\sqrt{N}}}\left(\phi\right)^{\otimes\lfloor rN\rfloor}\right)$ (400)